# Fast and accurate emulation of complex dynamic simulators Junoh Heo

### v1

@@ 3 Fast and exact emulation
In contrast to previously established findings, the prediction of GP with uncertain inputs in the one-step-ahead approach can be made computationally tractable by integrating the concept of Linked GP emulators. To be specific, the proposed approach predicts each state by linking GP outputs recursively, preserving accuracy without the need for sampling. Interestingly, the model structure depicted in Figure 1 can be interpreted as a special case within the emerging field of Deep GP (DGP) models (Damianou and Lawrence,, 2013), which have gained significant attention in the computer experiments literature for their hierarchical representation and enhanced flexibility (Sauer et al.,, 2023; Ming et al.,, 2023). Unlike canonical DGPs that rely on latent variable hierarchies, the proposed framework employs GP emulators fitted to each state variable, structured in a feed-forward manner reminiscent of deep learning models. This formulation facilitates the propagation of uncertainty across layers while maintaining analytical tractability, thereby enhancing predictive performance in dynamic system emulation. Building on this integration, we introduce a fast and accurate prediction algorithm for the one-step-ahead approach.
@@
Figure 1: An illustration of the self-linked GP emulator for a dynamic system governed by $d$ ODEs. Starting with the initial state $\mathbf{x}(t_{0})$, each emulator $\hat{f}_{m}$, $m=1,\ldots,d$ predicts the $m$-th component of the state $\hat{x}_{m}(t_{s})$ for $s=1,\ldots,T$. The predicted state $\hat{\mathbf{x}}(t_{s})$ is then fed forward as the input for the next time step $s+1$.
3.1 Algorithm without forcing inputs
For a nonlinear dynamic simulator $f$ governed by $d$ ODEs, consider GP emulators $\hat{f}_{m}$, $m=1,\ldots,d$ constructed by using an initial design $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, and corresponding outputs $\mathbf{y}_{m}=(x_{m}^{1}(t_{1}),\ldots,x_{m}^{n}(t_{1}))^{T}$ following the one-step-ahead approach. Instead of relying on MC approximations to compute (3) and (2.3), we derive analytical expressions for $\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ and $\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ under the squared exponential kernel given $\hat{\mathbf{x}}(t_{s-1})$ are as follows:
$\mu_{m}(\hat{\mathbf{x}}(t_{s}))$ $=\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]=\mathbb{E}\left[\hat{f}_{m}(\hat{\mathbf{x}}(t_{s-1}))\right]$
$=\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}+\sum^{n}_{i=1}r_{i}\prod_{l=1}^{d}\sqrt{\frac{\theta_{ml}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}\right)},$ (5)
@@ 3.1 Algorithm without forcing inputs
$\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s}))$ $=\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]=\mathbb{V}\left[\hat{f}_{m}(\hat{\mathbf{x}}(t_{s-1}))\right]$
$=\tau^{2}_{m}-\left(\mu_{m}(\hat{\mathbf{x}}(t_{s}))-\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}\right)^{2}+\left(\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{l=1}^{d}\zeta_{lik}\right),$ (6)
where $r_{i}=(\mathbf{K}^{-1}_{m}(\mathbf{y}_{m}-\mathbf{X}\beta_{m}))_{i}$, and
$\zeta_{lik}=\sqrt{\frac{\theta_{ml}}{\theta_{ml}+4\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}}\exp{\left(-\frac{\left(\frac{x^{i}_{l}(t_{0})+x^{k}_{l}(t_{0})}{2}-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\frac{\theta_{ml}}{2}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}-\frac{(x^{i}_{l}(t_{0})-x^{k}_{l}(t_{0}))^{2}}{2\theta_{ml}}\right)}.$
The closed-form expressions for the posterior predictive mean (3) and variance (2.3) under a Matérn kernel with smoothness parameters $\nu=1.5$ and $\nu=2.5$ can be straightforwardly derived following the proof in Ming and Guillas, (2021). Consistent with prior studies (Girard et al.,, 2002; Kyzyurova et al.,, 2018; Mohammadi et al.,, 2019; Ming and Guillas,, 2021), we adopt the moment matching method to approximate the posterior distribution as Gaussian, leveraging the mean and variance from (5) and (6). The moment matching approximation suits the structure of self-linked GP emulator, where outputs recursively serve as inputs, since the method aligns with the inherent Gaussian nature of GPs and ensures analytical tractability and recursive consistency.
Algorithm 1 Exact emulation of dynamic simulators without forcing inputs
@@
2:     Fit $\hat{f}_{m}$ using $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, and $\mathbf{y}_{m}=(x^{1}_{m}(t_{1}),\ldots,x^{n}_{m}(t_{1}))^{T}$
@@
4:Set $\hat{\mathbf{x}}(t_{0})\leftarrow\mathbf{x}(t_{0})$
5:for $s=0$ to $T-1$ do
@@
7:         Compute $\mu_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (5)
8:         Compute $\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (6)
@@
12:Return: $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$
By leveraging these closed-form expressions, computationally expensive MC approximations can be avoided, allowing for exact computations. The elimination of MC sampling significantly reduces computational costs, making the proposed approach particularly suitable for dynamical systems with many state variables that require extensive simulations. Moreover, the proposed framework provides exact predictive moments, ensuring both efficiency and accuracy. This is particularly advantageous for long-term iterative predictions, where even small numerical errors in early steps can accumulate over time, leading to significant deviations in later steps—a phenomenon often referred to as the butterfly effect. By directly computing the mean and variance, our approach mitigates this issue, enhancing robustness in sequential simulations. Furthermore, multi-step-ahead predictions can be efficiently obtained by recursively applying (5) and (6), reducing the need for costly recomputation at each iteration.
3.2 Algorithm with forcing input
Assume we have a known, $d_{w}$-dimensional forcing input $\mathbf{w}(t_{s})$ at time $t_{s}$ for $s=1,\ldots,T$. The dynamic simulator $f$ can be expressed as:
$x_{m}(t_{s+1})=f_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),\quad\text{for}\quad m=1,\ldots,d.$ (7)
Accordingly, the Markov property must be adjusted to account for forcing inputs. Instead of (2), the emulator follows:
$\hat{x}_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),$ (8)
where $m=1,\ldots,d$. The algorithm proposed in Section 3.1 can be naturally extended to incorporate forcing inputs by treating them as additional variables in the GP framework. This extension preserves the computational efficiency of the method while enabling accurate predictions for systems influenced by external inputs.
To account for forcing inputs, we modify the analytical expressions for $\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ and $\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ under the squared exponential kernel given $\hat{\mathbf{x}}(t_{s-1})$ are as follows:
$\mu_{m}(\hat{\mathbf{x}}(t_{s}))=\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}$
$+\sum^{n}_{i=1}r_{i}\prod_{j=1}^{d_{w}}\exp\left(-\frac{\left(w^{i}_{j}(t_{0})-w_{j}(t_{s})\right)^{2}}{\theta_{wj}}\right)\prod_{l=1}^{d}\frac{1}{\sqrt{1+2\frac{\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}{\theta_{ml}}}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mu_{l}(\hat{\mathbf{x}}(t_{s-1}))\right)^{2}}{\theta_{ml}+2\sigma^{2}_{l}(\hat{\mathbf{x}}(t_{s-1}))}\right)},$ (9)
and
$\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s}))=\tau^{2}_{m}-\left(\mu_{m}(\hat{\mathbf{x}}(t_{s}))-\begin{pmatrix}1&\hat{\mathbf{x}}(t_{s-1})\end{pmatrix}\beta_{m}\right)^{2}$
$+\left(\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{j=1}^{d_{w}}\exp{\left(-\frac{\left(w^{i}_{j}(t_{0})-(w_{j}(t_{s})\right)^{2}+\left(w^{k}_{j}(t_{0})-w_{j}(t_{s})\right)^{2}}{\theta_{wj}}\right)}\prod_{l=1}^{d}\zeta_{lik}\right),$ (10)
where $\theta_{wj}$ for $j=1,\ldots,d_{w}$ are the lengthscale hyperparameters associated with the forcing inputs.
Algorithm 2 Exact emulation of dynamic simulators with forcing inputs
1:for $m=1$ to $d$ do
2:     Fit $\hat{f}_{m}$ using $\mathbf{X}=(\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{n}(t_{0}))$, $\mathbf{W}=(\mathbf{w}^{1}(t_{0}),\ldots,\mathbf{w}^{n}(t_{0}))$, and $\mathbf{y}_{m}=(x^{1}_{m}(t_{1}),\ldots,x^{n}_{m}(t_{1}))^{T}$
3:end for
4:Set $\hat{\mathbf{x}}(t_{0})\leftarrow\mathbf{x}(t_{0})$
5:for $s=0$ to $T-1$ do
6:     for $m=1$ to $d$ do
7:         Compute $\mu_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (9) with $\mathbf{w}(t_{s})$
8:         Compute $\sigma^{2}_{m}(\hat{\mathbf{x}}(t_{s+1}))$ using (10) with $\mathbf{w}(t_{s})$
9:     end for
10:     Update $s\leftarrow s+1$
11:end for
12:Return: $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$
The additional term in these expressions corresponds to the kernel function $K(\mathbf{w}(t_{0}),\mathbf{w}(t_{s}))$, which arises because the forcing inputs $\mathbf{w}(t_{s})$ are treated as known rather than random. This assumption simplifies the model structure, as the forcing inputs remain unlinked variables in the emulator.
@@ 4 Numerical Studies
In this section, we conduct a series of numerical experiments to evaluate the performance of the proposed approach. Specifically, we compare the predictive performance of the proposed exact algorithm with an emulation method based on MC approximation, using $n_{MC}=1000$ samples. Both approaches employ the same GP emulator, using an anisotropic squared exponential kernel and a linear mean function $\alpha(\mathbf{x})=(1,\mathbf{x}^{T})\beta$.
All numerical studies are conducted using a maximin Latin hypercube design (Johnson et al.,, 1990; Morris and Mitchell,, 1995), with the number of initial samples set to $n=12d$, following the guidelines of Mohammadi et al., (2019), which builds on the design principles established in Jones et al., (1998) and Loeppky et al., (2009). The time step difference is set to $\Delta t=t_{s}-t_{s-1}=0.01$.
@@
To evaluate predictive performance, we compute predictions at all time points $t_{s}$, $s=1,\ldots,T$, and assess model accuracy using the root-mean-square error (RMSE) and continuous rank probability score (CRPS) (Gneiting and Raftery,, 2007) over 100 repetitions. The evaluation metrics are defined as follows:
RMSE $=\sqrt{\sum_{i=1}^{n}\frac{\left(y_{i}-\hat{y}_{i}\right)^{2}}{n}},$
CRPS $=\sigma\left[\frac{1}{\sqrt{\pi}}-2\phi\left(\frac{y-\mu}{\sigma}\right)-\frac{y-\mu}{\sigma}\left(2\Phi\left(\frac{y-\mu}{\sigma}\right)-1\right)\right],$
where $y$ represents the true function values, $\mu$ and $\sigma^{2}$ denote the posterior predictive mean and variance, respectively. $\phi$ and $\Phi$ are the pdf and the cdf of a standard Gaussian distribution. These metrics quantify the degree of discrepancy between the true function and the predictive model, with lower RMSE and CRPS values indicating better model accuracy. Furthermore, we evaluate computational efficiency by comparing computation times across different emulation approaches.
@@ 4.1 Lotka-Volterra model
where $r_{G}>0$ represents the intrinsic growth rate of the prey in the absence of predators, $K$ is the prey carrying capacity, $r_{I}>0$ is the predation rate coefficient, $k_{AE}>0$ is the assimilation efficiency of the predator, and $r_{M}>0$ denotes the natural mortality rate of the predator in the absence of prey. For this study, we consider a simulation over $t\in[0,30]$, with prespecified parameter values $r_{G}=1.5$, $K=10$, $r_{I}=1.5$, $k_{AE}=1$, and $r_{M}=2$. Initial samples are drawn from $[0,5]^{2}$. The objective is to predict the system’s dynamics given the initial state $\mathbf{x}(t_{0})=(P(t_{0})=1,C(t_{0})=2)$.
@@
Figure 2: RMSE (first and second plots), CRPS (third and fourth plots) of emulation algorithms to $P$ and $C$ and computation time (fifth plots) across 100 repetitions for Lotka-Volterra model.
Figure 3: Lotka-Volterra model (black solid line) and emulation results of prediction (green dashed line) via each algorithm.
@@
Figure 2 presents the emulation results over 100 repetitions for RMSE and CRPS, while Figure 3 illustrates the predicted trajectories for state variables, $P$ and $C$, respectively. As expected, both algorithms produce nearly identical predictions, with a maximum difference of only 0.000024. While both methods achieve high predictive accuracy, their computational costs differ considerably. On average, the exact algorithm requires only 0.40603 seconds per repetition, whereas the MC approach takes 90.13912 seconds. This comparison highlights the advantage of the proposed framework in significantly reducing computational cost while maintaining accuracy in dynamic system emulation. Moreover, the computational gap would increase further if the MC approach employed a larger number of samples, or if the simulation involved many time steps or a longer time horizon.
Figure 4: Boxplots of computation time for the Exact and Monte Carlo methods as the number of design points varies from 5 to 50 in increments of 5 across 100 repetitions for Lotka-Volterra model.
Figure 4 illustrates the computation time for emulation as the number of samples varies from 5 to 50 per repetition. With a precomputed $\mathbf{K}^{-1}$, the computational complexity of computing the predictive mean and variance at a single input location using MC-based emulation is $\mathcal{O}(n^{2}n_{MC})$. On the other hand, under the same conditions, the exact algorithm requires only $\mathcal{O}(n^{2}d)$ complexity. Since $n_{MC}$ is typically in the range of several thousand or more, the proposed algorithm offers substantial computational savings.

### v2

@@ 3 Fast and exact emulation
The one-step-ahead emulation framework can be interpreted as a probabilistic analog of the Euler method (Biswas et al.,, 2013) for solving ODEs. The classical Euler scheme approximates the evolution of a dynamical system by discretizing time and iteratively applying
$\mathbf{x}(t_{s+1})=\mathbf{x}(t_{s})+\Delta t\cdot h(\mathbf{x}(t_{s})),$
where $h$ denotes the function governing the dynamics of the system. In our setting, this transition is modeled as a data-driven one-step mapping $x(t_{s})\to x(t_{s+1})$, learned using GP emulators. Importantly, unlike previously established findings (Girard et al.,, 2002; Mohammadi et al.,, 2019) that rely on MC approximation, our method analytically derives the predictive mean and variance using linked GPs, enabling exact and sampling-free uncertainty propagation at each time step. This allows the GP-based flow map to serve as a probabilistic surrogate for the unknown vector field $h$, yielding a data-driven and uncertainty-aware generalization of the Euler scheme.
Building on this foundation, we integrate the concept of linked GP emulators to propagate predictions with uncertain inputs over time. Specifically, our method recursively links GP outputs to emulate each state variable, preserving accuracy while maintaining computational efficiency. The feed-forward architecture of the proposed framework, illustrated in Figure 1, forms a sequentially layered structure where each predicted state is linked to the previous one. Interestingly, this structure can be interpreted as a special case within the emerging field of deep GP (DGP) models (Damianou and Lawrence,, 2013), which have gained significant attention in the computer experiments literature for their hierarchical representation and enhanced flexibility (Ko and Kim,, 2022; Sauer et al.,, 2023; Ming et al.,, 2023). However, unlike canonical DGPs that involve latent intermediate layers and require approximate inference, our approach replaces those latent layers with a series of GP emulators, linked recursively to propagate states in a feed-forward manner reminiscent of deep learning models.
@@
Figure 1: An illustration of the self-linked GP emulator for a dynamic system governed by $d$ ODEs. Starting with the initial state $\mathbf{x}(t_{0})$, each emulator $\hat{f}_{m}$, $m=1,\ldots,d$ predicts the $m$-th component of the state $x_{m}(t_{s})$ for $s=1,\ldots,T$. The predicted state $\mathbf{x}(t_{s})$ is then fed forward as the input for the next time step $s+1$.
We now detail the proposed exact and sampling-free, recursive emulation procedure for the dynamic simulator $f$ introduced in Section 2.3. As specified in equation (2), the simulator evolves via the one-step transition $\mathbf{x}(t_{s+1})=(\hat{f}_{1}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),\ldots,\hat{f}_{d}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})))^{\top}$, where the forcing input $\mathbf{w}(t_{s})$ may be omitted in autonomous systems. To emulate this mapping, we construct $d$ independent GP models $\hat{f}_{m}$, one for each state variable $m=1,\ldots,d$, using the one-step-ahead responses $\mathbf{y}_{m}=(x_{m}^{1}(t_{1}),\ldots,x_{m}^{n}(t_{1}))$. Depending on the presence of forcing inputs, the emulator takes the form either $x_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))$ or $x_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}))$.
Predictions are obtained recursively using the following closed-form expressions for the posterior predictive mean in (3) and variance in (2.3) under a squared exponential kernel, originally derived by Kyzyurova et al., (2018) and adapted here for the one-step-ahead emulation setting:
$\mathbb{E}(x_{m}(t_{s+1}))=\begin{pmatrix}1&\mathbf{x}(t_{s})\end{pmatrix}\beta_{m}$
$\quad+\sum^{n}_{i=1}r_{i}\prod_{j=1}^{p}\exp\left(-\frac{\left(w^{i}_{j}(t_{0})-w_{j}(t_{{s+1}})\right)^{2}}{\theta_{wj}}\right)\prod_{l=1}^{d}\sqrt{\frac{\theta_{ml}}{\theta_{ml}+2\mathbb{V}(x_{l}(t_{s}))}}\exp{\left(-\frac{\left(x^{i}_{l}(t_{0})-\mathbb{E}(x_{l}(t_{s}))\right)^{2}}{\theta_{ml}+2\mathbb{V}(x_{l}(t_{s}))}\right)},$ (5)
@@
$\mathbb{V}(x_{m}(t_{s+1}))=\tau^{2}_{m}-\left(\mathbb{E}(x_{m}(t_{s+1}))-\begin{pmatrix}1&\mathbf{x}(t_{s})\end{pmatrix}\beta_{m}\right)^{2}$
$\quad+\sum_{i,k=1}^{n}\left(r_{i}r_{k}-\tau^{2}_{m}(\mathbf{K}^{-1}_{m})_{ik}\right)\prod_{j=1}^{p}\exp{\left(-\frac{\left(w^{i}_{j}(t_{0})-w_{j}(t_{s+1})\right)^{2}+\left(w^{k}_{j}(t_{0})-w_{j}(t_{s+1})\right)^{2}}{\theta_{wj}}\right)}\prod_{l=1}^{d}\zeta_{lik},$ (6)
where $r_{i}=(\mathbf{K}^{-1}_{m}(\mathbf{y}_{m}-\begin{pmatrix}1&\mathbf{X}\end{pmatrix}\beta_{m}))_{i}$ and
$\zeta_{lik}=\sqrt{\frac{\theta_{ml}}{\theta_{ml}+4\mathbb{V}(x_{l}(t_{s}))}}\exp{\left(-\frac{\left(\frac{x^{i}_{l}(t_{0})+x^{k}_{l}(t_{0})}{2}-\mathbb{E}(x_{l}(t_{s}))\right)^{2}}{\frac{\theta_{ml}}{2}+2\mathbb{V}(x_{l}(t_{s}))}-\frac{(x^{i}_{l}(t_{0})-x^{k}_{l}(t_{0}))^{2}}{2\theta_{ml}}\right)}.$
Here, $\theta_{ml}$ and $\theta_{wj}$ are the lengthscale hyperparameters associated with the state and forcing input dimensions, respectively.
Algorithm 1 Exact emulation of dynamic simulators
@@
2:  Fit $\hat{f}_{m}$ on training data $(\mathbf{X},\mathbf{W},\mathbf{y}_{m})$
@@
4:Set $\mathbf{x}(t_{0})\leftarrow\mathbf{x}(t_{0})$
5:for $s=0$ to T1T-1 do
@@
7:   Compute $\mathbb{E}(x_{m}(t_{s+1}))$ using (5)
8:   Compute $\mathbb{V}(x_{m}(t_{s+1}))$ using (6)
@@
12:Return: $\mathbf{x}(t_{s})$ for $s=1,\ldots,T$
The forcing inputs are treated as known and contribute deterministically through their kernel terms, without introducing additional uncertainty. When forcing inputs are absent, the expressions simplify by setting $p=0$. The closed-form expressions for the posterior predictive mean and variance under a Matérn kernel can be straightforwardly derived following the development in Ming and Guillas, (2021).
Consistent with prior studies (Girard et al.,, 2002; Kyzyurova et al.,, 2018; Mohammadi et al.,, 2019; Ming and Guillas,, 2021), we adopt the moment matching method to approximate the posterior distribution as Gaussian, leveraging the analytic mean and variance in (5) and (6). This approach is particularly well-suited to the self-coupled structure of the proposed framework, where the prediction at each step serve as input to the next time step.
By computing the mean and variance analytically, the proposed sampling-free method achieves exact moment propagation, avoiding the accumulation of sampling error. This is especially beneficial for predictions over long time horizon, where even small numerical inaccuracies from early steps can compound over time, leading to significant degradations in later stages—a phenomenon commonly referred to as the butterfly effect. Moreover, the feed-forward structure supports efficient multi-step-ahead predictions without repeated recomputation of previous states, providing further computational gains.
@@ 4 Numerical Studies
In this section, we conduct a series of numerical experiments to evaluate the performance of the proposed approach. Specifically, we compare the predictive performance of the proposed exact algorithm with an emulation method based on MC approximation, considering three sample sizes $(n_{MC}=10,100,1000)$. Importantly, both approaches adopt the moment-matching method, and employ the same underlying GP emulator, constructed using an anisotropic squared exponential kernel and a linear mean function $\alpha(\mathbf{x})=(1,\mathbf{x})\beta$. The difference between the two methods lies solely in how predictive moments are computed as described in Section 2: the MC-based approach relies on sampling, whereas the proposed exact method computes them analytically.
All numerical experiments are conducted using a maximin Latin hypercube design (Johnson et al.,, 1990; Morris and Mitchell,, 1995), with the number of initial samples set to $n=10\times d$, following the guidelines of Jones et al., (1998) and Loeppky et al., (2009). The time step difference is set to $\Delta t=t_{s}-t_{s-1}=0.01$.
@@
Predictive performance is evaluated across all time points $t_{s}$, $s=1,\ldots,T$, using two standard metrics: the root-mean-square error (RMSE) and the mean absolute error (MAE), averaged over 100 repetitions. The evaluation metrics are defined as follows:
RMSE $=\sqrt{\frac{1}{T}\sum_{s=1}^{T}\left(x_{m}(t_{s})-\mathbb{E}(x_{m}(t_{s}))\right)^{2}},$
MAE $=\frac{1}{T}\sum_{s=1}^{T}|x_{m}(t_{s})-\mathbb{E}(x_{m}(t_{s}))|.$
These metrics measure the discrepancy between the true function and the predictive model, with lower RMSE and MAE values indicating higher accuracy. Furthermore, computational efficiency is assessed by comparing computation times across different emulation approaches.
In addition to prediction accuracy, we assess how predictive uncertainty accumulates over time. Specifically, we determine the number of steps for which predictions remain reliable and the point at which uncertainty increases substantially. For this purpose, we apply change point detection (Killick et al.,, 2012) to the predictive standard deviation using the cpt.mean function in the changepoint package (Killick and Eckley,, 2014). The comparison is performed only between the exact method and the MC approach with 1000 samples, as the latter provides the most accurate MC-based predictions and yields more stable results by reducing variability from sampling. This analysis addresses the practical question of how far ahead the emulator can provide useful predictions before the uncertainty becomes large enough to degrade performance.
@@ 4.1 Lotka-Volterra model
where $r_{G}>0$ represents the intrinsic growth rate of the prey in the absence of predators, $K$ is the prey carrying capacity, $r_{I}>0$ is the predation rate coefficient, $k_{AE}>0$ is the assimilation efficiency of the predator, and $r_{M}>0$ denotes the natural mortality rate of the predator in the absence of prey. For this study, we consider a simulation over $t\in[0,30]$, with prespecified parameter values $r_{G}=1.5$, $K=10$, $r_{I}=1.5$, $k_{AE}=1$, and $r_{M}=2$. Initial samples are drawn from $0,5]^{2}$. The objective is to predict the system’s dynamics given the initial state $\mathbf{x}(t_{0})=(P(t_{0})=1,C(t_{0})=2)$.
@@
Figure 2: RMSE (first and second panels), MAE (third and fourth panels) of emulation algorithms to $P$ and $C$ and computation time (fifth panel) across 100 repetitions for the Lotka-Volterra model.
@@
Figure 2 presents the emulation results over 100 repetitions for RMSE and MAE. As expected, the MC approach produces suboptimal performance, although larger sample sizes yield results closer to those of the exact method. Compared with the exact method, the MC approach with 10 samples yields RMSE and MAE values that are 6–8% larger, with 100 samples yields values that are 4–6% larger, and with 1000 samples yields values that are less than 1% larger. Moreover, their computational costs differ considerably. On average, the exact algorithm requires only 0.45031 seconds per repetition, whereas the MC approach takes 0.48785 seconds with 10 samples, 1.70622 seconds with 100 samples, and 101.37841 seconds with 1000 samples. This comparison highlights the advantage of the proposed framework in significantly reducing computational cost while maintaining accuracy in dynamic system emulation. As a result, the computational gap would increase further if the MC approach employed a larger number of samples, or if the simulation involved many time steps or a longer time horizon.
Figure 3: Lotka-Volterra model (black solid line) and emulation results of prediction (green dashed line) via each algorithm with detected change points (red dotted vertical line)
Figure 3 illustrates the predicted trajectories for state variables, $P$ and $C$, along with the change point of predictive uncertainty. Because the one-step-ahead approach is effective enough for emulating the Lotka-Volterra model, both methods accurately predict the simulator outputs with very small uncertainties of approximately $10^{-4}$ for the exact method and $10^{-2}$ for MC method. The Exact method does not detect any change points within the given time steps, whereas the MC method detects changes in uncertainties at $t=5.53$ for $P$ and $t=3.33$ for $C$.
4.2 Duffing oscillator
The Duffing oscillator (Duffing,, 1918; Ueda,, 1991) is a canonical model in nonlinear dynamics that describes the behavior of a damped oscillator subject to a nonlinear restoring force. It captures the dynamics of a second-order system defined in terms of the displacement variable $x(t)$ and its velocity $v(t)$, governed by the following system of differential equations:
$\begin{cases}\frac{dx}{dt}&=v,\\ \frac{dv}{dt}&=-\delta v-\alpha x-\beta x^{3},\end{cases}$
where $\delta>0$ is the damping coefficient, $\alpha$ controls the linear stiffness, and $\beta$ governs the strength of the nonlinear restoring force. In this study, we simulate the system over $t\in[0,30]$ with parameter values $\delta=0.1$, $\alpha=1$, and $\beta=1$. The initial state is set as $\mathbf{x}(t_{0})=(x(t_{0})=1,v(t_{0})=0)$, and training inputs are sampled from the domain $1,1]^{2}$.